Please help me solve this equation Two consecutive numbersare such that the greater added thrice the smaller makes a total of 53
Answer by JulietG(1812)

B = A + 1 (consecutive)
B + 3A = 53
Substitute the value that we know for B
B + 3A = 53
(A+1) + 3A = 53
4A + 1 = 53
subtract 1 from each side
4A = 52
divide by 4
A = 13
.
If A is 13, then B is 14.
Let's make sure it works.
14 + (3*13) = 53
14 + 39 does indeed equal 53
success!